if f(x)=(ax^2+b)^3 then the function g such that f(g(x))=g(f(x)) is given by

Grade:12th pass

1 Answers

Shubham Dobliyal
6 years ago
According to me, there are two ways:One can be that g(x) is inverse of f(x). Then f(g(x)) =g(f(x)) =x.... For that g(x) =((x^1/3-b)/a)^1/2Other solution can be when g(x) =f(x) =(ax^2+b)^3Hope you understand